OpenStudy (anonymous):

solve -2x - 2 (-7x + 7) = 82

OpenStudy (photon336):

okay so you see that -2 on the outside of the ( ) ?

OpenStudy (anonymous):

yes

OpenStudy (photon336):

we have to distribute that -2 in the inside so that means we multiply -2 by everything in the ( ) so for this part here's what we have \[-2(-7x+7) = 14x-14\]

OpenStudy (anonymous):

ok I'm following

OpenStudy (photon336):

now so the next step is here is what we have -2x-14x-14 = 82 we've got to combine like terms, how do you think we would do this?

OpenStudy (anonymous):

-16x - 14 = 82

OpenStudy (photon336):

yes now we can easily isolate x. what's our next step?

OpenStudy (anonymous):

add 14 to both sides

OpenStudy (photon336):

yes

OpenStudy (photon336):

so now we've got -16x = 96

OpenStudy (anonymous):

x = -6

OpenStudy (photon336):

and that's all it is. the most important thing was to distribute the negative sign outside of the ( )

OpenStudy (anonymous):

thanks! i have another question about something else

OpenStudy (photon336):

ok

OpenStudy (anonymous):

how do you put square root in a calculator?

OpenStudy (photon336):

well you do two things: you can take your number and put ( ) around it and raise it to the 1/2 power (n)^{0.5}

OpenStudy (photon336):

or you can press the sqrt button

OpenStudy (photon336):

\[\sqrt{n} = (n)^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

i'm trying to find that on my calculator

OpenStudy (anonymous):

the sqrt button

OpenStudy (anonymous):

i don't see it

OpenStudy (anonymous):

i have a TI-84 Plus

OpenStudy (photon336):

okay so you see this button \[x^{2}\]

OpenStudy (anonymous):

yes

OpenStudy (photon336):

press the second button then \[x^{2}\]

OpenStudy (anonymous):

ok

OpenStudy (photon336):

\[second \rightarrow x^{2}\]

OpenStudy (anonymous):

omg how did you know that? thanks so much!!!

OpenStudy (photon336):

When I was in HS i used a T183

OpenStudy (photon336):

same thing but less advanced

OpenStudy (anonymous):

omg that was so helpful. so how do i approximate the square root of 13 to the nearest tenth

OpenStudy (photon336):

that i'm not sure about

OpenStudy (photon336):

usually I just round it in my head when the answer comes out

OpenStudy (anonymous):

oh ok